• The electrical load in a small workshop consists of 14 lamps, each rated at 240volt, 60watt and 3 electric fire, each rated at 240, 1kw. What is the effective resistance of the load
1
Expert's answer
2021-10-18T03:45:02-0400

Resistance of a one lamp

"R_1=\frac{V_1^2}{P_1}=\frac{240^2}{60}=960\:\Omega"

Resistance of a one electric fire

"R_2=\frac{V_2^2}{P_2}=\frac{240^2}{1000}=57.6\:\Omega"

The net resistance

"R=\frac{1}{\frac{14}{R_1}+\frac{3}{R_2}}=\frac{1}{\frac{14}{960}+\frac{3}{57.6}}=15\:\Omega"